# XMAGNET : Kinetic, Thermal and Magnetic AGN Feedback in Massive Galaxies at Halo Masses $\sim 10^{13.5}$ M

### v1

@@ Abstract
Galaxy : active, Galaxies : intergalactic medium, galaxies : evolution, quasars : supermassive black holes, galaxies : nuclei, galaxies : magnetic fields, galaxies : jets, galaxies : star formation
@@ 1 Introduction
Galaxies can be supplied cold gas for star formation in multiple ways including cold streams (Kereš et al. 2005), stellar mass loss (Mathews & Brighenti 2003; McDonald et al. 2018), and in-situ cooling of the circumgalactic medium (CGM) due to thermal instability (Field 1965; White & Frenk 1991). The cold gas being fed through cold streams encounters hot CGM, which likely prevents it from reaching the core and thus also from fueling star formation in quiescent galaxies (Heitsch & Putman 2009). The gas being shed by dying stars provides another channel to supply gas for star formation. Energetics show that supernova explosions are capable of driving the stellar ejecta from the galaxy to the CGM, thus limiting star formation (Voit et al. 2015b). However, if the CGM pressure is large, supernovae cannot do so leaving that gas available for star formation within the galaxy (Voit et al. 2020). The cooling time ($t_{\rm cool}$) of the CGM surrounding massive galaxies is of the order of 100s Myr within the central $r<30$ kpc, much shorter than the galaxy’s lifetime (Hogan et al. 2017). As such, the CGM is expected to radiatively cool due to thermal instability and feed the cold gas into the galaxy, fueling star formation. If left uninterrupted, CGM cooling would become catastrophic, making massive galaxies appear blue due to very high star formation rates (Fabian 1994). However, observations of massive elliptical galaxies show them to be much more quiescent with respect to star formation (Combes et al. 2007), necessitating the need for some feedback mechanism to quench star formation.
@@
Grete et al. 2025 introduced the XMAGNET [^1] See https://xmagnet-simulations.github.io for further material and videos. (eXascale simulations of Magnetized AGN feedback focusing on Energetics and Turbulence) suite of simulations designed to explore the role of magnetised AGN feedback in massive galaxies, groups, and clusters. It discussed the role of magnetised and kinetic AGN feedback mechanisms in Perseus-type cool-core galaxy clusters ($M_{200}\sim 6.6\times 10^{15}\,M_{\odot}$). This paper focuses on the role of AGN feedback in lower mass galactic halos with $M_{200}\sim 10^{13.5}\,M_{\odot}$, in which AGN feedback using the same model couples with CGM pressure and Type Ia supernovae (SNIa) heating in the presence of magnetic fields.
@@ 2 Methods
Runs M300 c200 R200 $_{0}^{\dagger} K$$_{100}^{\dagger}$ $\alpha_{K}^{\dagger}$ MBH M $r_{H}$ B field Thermal
@@ 2.1 Simulation Grid
The simulation is set up in a (6.4 Mpc)3 cubic box, covered by $512^{3}$ cells in the static root grid. We enforce 7 levels of refinement within $([-25,25]\mathrm{kpc})^{3}$ (where the root grid is the 0th level). Thus, the central (50 kpc)3 of the setup domain is covered with a uniform grid of $512^{3}$ cells with cell side length of $\Delta x\approx 100$ pc.
@@ 2.2 Gravitational Potential
Table 1 lists the runs discussed in this paper along with the parameters used for initial setup. In the SPG and SPG with cooler core simulations a central galaxy with stellar mass $M_{*}=2\times 10^{11}\,M_{\odot}$ and central velocity dispersion $\sigma_{v}\approx 280\,{\rm km\,s^{-1}}$ is embedded in a halo of mass $M_{\rm halo}=10^{13.6}\,M_{\odot}$ ($c_{200}=7.5$) and has a central black hole mass $M_{\rm BH}=2.6\times 10^{9}\,M_{\odot}$. In the MPG simulations a central galaxy with stellar mass $M_{*}=1.2\times 10^{11}\,M_{\odot}$ and central velocity dispersion $\sigma_{v}\approx 230\,{\rm km\,s^{-1}}$ is embedded in a halo of mass $M_{\rm halo}=10^{13.6}\,M_{\odot}$ ($c_{200}=7.5$) and has a central black hole mass $M_{\rm BH}=4.6\times 10^{8}\,M_{\odot}$.
@@ 2.3 Initial Baryon Profile
As stated in Section 2.2 we initialize the baryon profile in hydrostatic equilibrium within the gravitational potential well given by the NFW+Hernquist+SMBH profile. Figure 1 shows the initial radial electron density (top panel), entropy (middle panel) and cooling time ($t_{\rm cool}$; lower panel) profile for the single phase galaxy (SPG; blue line), multiphase galaxy (MPG; red line) and galaxy with SPG potential with shallower entropy and higher CGM density and pressure (SPG-cool core; black line). For the SPG halo we have chosen NGC4472 as the analogue, and for the MPG halo NGC5044 has been chosen as the analogue. The SPG with cooler core has been initialised to study the interplay between AGN feedback, SNIa-driven sweeping of gas out of the interstellar medium (and thus supression of star formation), and CGM pressure in a situation where the central stellar velocity dispersion is $\sigma_{v}>240$ km s-1 (like the SPG case) but also has a much higher CGM pressure (like the MPG case) to start with. This will allow us to test whether AGN+SNIa feedback alone is able to reconfigure the CGM entropy profile to be similar to the SPG-type system – in other words, if it can be reconfigured by feedback alone to have the lower CGM density and pressure that observations tell us are appropriate for the central potential. (If this is not the case, this likely implies that either the AGN feedback has to be active during the halo assembly stage to keep the CGM pressure low or a merger between comparably massive cosmological halos is needed to reconfigure the CGM.) Figure 1 shows that the baryon profile for the SPG with cooler core is closer to that of the multiphase galaxy with higher CGM pressure and cooling time, $t_{\rm cool}\lesssim 0.5$ Gyr within the central ($r<10$ kpc). The parameter of the runs are listed in table 1.
@@ 3.2.1 Temporal evolution
Figure 5 shows the evolution of the total cold gas mass ($T<10^{5}$ K), jet power (Pjet), total stellar mass, and X-ray luminosity ($L_{X}$) within $r=50$ kpc for the 0.570.5-7 keV gas for the SPG-MHD (left panel) and SPG-hydro (right panel) runs. These plots show a tight correlation between jet power ($P_{\rm jet}$) and X-ray luminosity ($L_{X}$; 0.5 keV<T<7<T<7keV),as every AGN outburst is preceded by a rise in $L_{X}$. For both runs, cold gas mass remains $\lesssim 10^{7}\ {\rm M}_{\odot}$ resulting in AGN activity with peak power (${\rm P_{jet}}\sim 10^{43}$ erg s-1). The jet duty cycle is typically $\sim 100$ Myr for both the runs with no star formation during their evolution until at least $t=2$ Gyr. The cold gas formation remains centrally concentrated within $r=1$ kpc with no extended cold gas filament formation (see Figure 17).
@@ 4.5 Comparison with other works
The analysis presented in Voit et al. (2015a); Voit et al. (2020) has motivated multiple simulations exploring the interplay between radiative cooling, stellar feedback and AGN feedback in the evolution of massive elliptical galaxies. Wang et al. (2019) performed hydrodynamic simulations focused on galaxies similar to our single phase and multiphase galaxies. Similar to Prasad et al. (2020), they also find that in single phase galaxy AGN feedback is able to maintain a steady hot CGM with a small amount of centrally-concentrated cold gas. A similar mass multiphase galaxy forms a large amount of extended cold gas. Our simulations are largely in line with these findings for the single phase and multiphase galaxies despite having very different numerics and physics modules. Our results deviate from Prasad et al. (2020) in the central $r\simeq 3$ kpc, however, as we see a flattening of the entropy profile for both SPG-hydro and MPG-hydro simulations. We speculate that this is due to the strong thermal feedback in our simulations within $r\simeq 1$ kpc. Our MHD runs also show similar behaviour to the hydro runs with partial thermal AGN feedback in the radial profile. The spherical heat injection by AGN feedback in addition to the thermal feedback due to Type Ia supernova feedback leads to overheating within the central $r\simeq 3$ kpc. Recent work by Mohapatra et al. (2025) suggests that modeling the Type Ia supernova feedback via discrete explosions rather than uniformly distributed heating helps to maintain the low central entropy within $r\simeq 5$ kpc.
@@ Acknowledgements
Data available on request.

### v2

@@ Abstract
Galaxies, Cooling Flow, Super Massive Black Hole, AGN Feedback, Stellar Feedback, Circumgalactic Medium
@@ 1 Introduction
Galaxies can be supplied cold gas for star formation in multiple ways including cold streams (Kereš et al. 2005), stellar mass loss (Mathews & Brighenti 2003), and in-situ cooling of the circumgalactic medium (CGM) due to thermal instability (Field 1965; White & Frenk 1991). The cold gas being fed through cold streams encounters hot CGM, which likely prevents it from reaching the core and thus also from fueling star formation in quiescent galaxies (Heitsch & Putman 2009). The gas being shed by dying stars provides another channel to supply gas for star formation. Energetics show that supernova explosions are capable of driving the stellar ejecta from the galaxy to the CGM, thus limiting star formation (Voit et al. 2015b). However, if the CGM pressure is large, supernovae cannot do so leaving that gas available for star formation within the galaxy (Voit et al. 2020). The cooling time ($t_{\rm cool}$) of the CGM surrounding massive galaxies is of the order of 100s Myr within the central $r<30$ kpc, much shorter than the galaxy’s lifetime (Hogan et al. 2017). As such, the CGM is expected to radiatively cool due to thermal instability and feed the cold gas into the galaxy, fueling star formation. If left uninterrupted, CGM cooling would become catastrophic, making massive galaxies appear blue due to very high star formation rates (Fabian 1994). However, observations of massive elliptical galaxies show them to be much more quiescent with respect to star formation (Combes et al. 2007), necessitating the need for some feedback mechanism to quench star formation.
@@
Grete et al. 2025 introduced the XMAGNET [^1] See https://xmagnet-simulations.github.io for further material and videos. (eXascale simulations of Magnetized AGN feedback focusing on Energetics and Turbulence) suite of simulations designed to explore the role of magnetised AGN feedback in massive galaxies, groups, and clusters. It discussed the role of magnetised and kinetic AGN feedback mechanisms in Perseus-type cool-core galaxy clusters ($M_{200}\sim 6.6\times 10^{15}\,M_{\odot}$). In this work, we focus on the role of AGN feedback in lower mass galactic halos with $M_{200}\sim 10^{13.5}\,M_{\odot}$, in which AGN feedback using the same model couples with CGM pressure and Type Ia supernovae (SNIa) heating in the presence of magnetic fields. The paper proceeds as follows : Section 2 introduces the numerical and analysis methods for our simulations, Section 3 presents the main results of our simulations, Section 4 looks more closely at how AGN feedback couples with CGM, its larger implications and compares this work with similar studies, and finally in Section 5, we present the conclusions based on our findings.
@@ 2 Methods
[…]
@@
  • $\beta$ - Parameters for the BCG’s Hernquist profile : $\Phi_{\rm BCG}=\frac{GM}{r+r_{H}}$
  • SMBH mass, $M_{\rm BH}$, is taken from Kormendy & Ho (2013) for SPG and from David et al. (2009) for the MPG halo.
@@ 2.1 Simulation Grid
The simulation is set up in a (6.4 Mpc)3 cubic box, covered by $512^{3}$ cells in the static root grid. We enforce 7 levels of refinement within $([-25,25]\mathrm{kpc})^{3}$ (where the root grid is the 0th level). Thus, the central (50 kpc)3 of the setup domain is covered with a uniform grid of $512^{3}$ cells with cell side length of $\Delta x\approx 100$ pc. Outside of this central region the resolution is progressively coarsened by factors of 2 (with intervening regions of constant resolution either 256 or 512 cells wide) to ensure that the transition in resolution is smooth and all phenomena of interest are sufficiently resolved.
@@ 2.2 Gravitational Potential
Table 1 lists the runs discussed in this paper along with the parameters used for initial setup. In the SPG (halo with deeper central potential, $\sigma_{v}>240$ km s-1, and lower CGM pressure) and SPG with cooler core (Halo with deeper central potential, $\sigma_{v}>240$ km s-1, and higher CGM pressure) simulations, a central galaxy with stellar mass $M_{*}=2\times 10^{11}\,M_{\odot}$ and central velocity dispersion $\sigma_{v}\approx 280\,{\rm km\,s^{-1}}$ is embedded in a halo of mass $M_{\rm halo}=10^{13.6}\,M_{\odot}$ ($c_{200}=7.5$) and has a central black hole mass $M_{\rm BH}=2.6\times 10^{9}\,M_{\odot}$. In the MPG simulations (halo with shallower central potential, $\sigma_{v}<240$ km s-1, and higher CGM pressure), a central galaxy with stellar mass $M_{*}=1.2\times 10^{11}\,M_{\odot}$ and central velocity dispersion $\sigma_{v}\approx 230\,{\rm km\,s^{-1}}$ is embedded in a halo of mass $M_{\rm halo}=10^{13.6}\,M_{\odot}$ ($c_{200}=7.5$) and has a central black hole mass $M_{\rm BH}=4.6\times 10^{8}\,M_{\odot}$.
@@ 2.3 Initial Baryon Profile
As stated in Section 2.2 we initialize the baryon profile in hydrostatic equilibrium within the gravitational potential well given by the NFW+Hernquist+SMBH profile. Figure 1 shows the initial radial electron density (top panel), entropy (middle panel) and cooling time ($t_{\rm cool}$; lower panel) profile for the single phase galaxy (SPG; blue line), multiphase galaxy (MPG; red line) and galaxy with SPG potential with shallower entropy and higher CGM density and pressure (SPG-cool core; black line). For the SPG halo we have chosen NGC4472 as the analogue (Werner et al. 2014), and for the MPG halo NGC5044 has been chosen as the analogue (Werner et al. 2014). The SPG with cooler core has been initialised to study the interplay between AGN feedback, SNIa-driven sweeping of gas out of the interstellar medium (and thus supression of star formation), and CGM pressure in a situation where the central stellar velocity dispersion is $\sigma_{v}>240$ km s-1 (like the SPG case) but also has a much higher CGM pressure (like the MPG case) to start with. This will allow us to test whether AGN+SNIa feedback alone is able to reconfigure the CGM entropy profile to be similar to the SPG-type system – in other words, if it can be reconfigured by feedback alone to have the lower CGM density and pressure that observations tell us are appropriate for the central potential. (If this is not the case, this likely implies that either the AGN feedback has to be active during the halo assembly stage to keep the CGM pressure low or a merger between comparably massive cosmological halos is needed to reconfigure the CGM.) Figure 1 shows that the baryon profile for the SPG with cooler core is closer to that of the multiphase galaxy with higher CGM pressure and cooling time, $t_{\rm cool}\lesssim 0.5$ Gyr within the central ($r<10$ kpc). The parameter of the runs are listed in table 1.
@@ 3.2.1 Temporal evolution
Figure 5 shows the evolution of the total cold gas mass ($T<10^{5}$ K), jet power (Pjet), total stellar mass, and X-ray luminosity ($L_{X}$) within $r=50$ kpc for the 0.570.5-7 keV gas for the SPG-MHD (left panel) and SPG-hydro (right panel) runs. These plots show a tight correlation between jet power ($P_{\rm jet}$) and X-ray luminosity ($L_{X}$; 0.5 keV<T<7<T<7keV),as every AGN outburst is preceded by a rise in $L_{X}$. For both runs, cold gas mass remains $\lesssim 10^{7}\ {\rm M}_{\odot}$ resulting in AGN activity with peak power (${\rm P_{jet}}\sim 10^{43}$ erg s-1). The jet duty cycle is typically $\sim 100$ Myr for both the runs with no star formation during their evolution until at least $t=2$ Gyr. The cold gas formation remains centrally concentrated within $r=1$ kpc with no extended cold gas filament formation (see Figure 17). As such, the heating-cooling cycle remains very tightly coupled unlike the MPG runs, where extended cold gas formation leads to much longer jet events with an order of magnitude higher jet power.
@@ 4.5 Comparison with other works
The analysis presented in Voit et al. (2015a); Voit et al. (2020) has motivated multiple simulations exploring the interplay between radiative cooling, stellar feedback and AGN feedback in the evolution of massive elliptical galaxies. Wang et al. (2019) performed hydrodynamic simulations focused on galaxies similar to our single phase and multiphase galaxies. Similar to Prasad et al. (2020), they also find that in single phase galaxy AGN feedback is able to maintain a steady hot CGM with a small amount of centrally-concentrated cold gas. A similar mass multiphase galaxy forms a large amount of extended cold gas. Our simulations are largely in line with these findings for the single phase and multiphase galaxies despite having very different numerics and physics modules. Our results deviate from Prasad et al. (2020) in the central $r\simeq 3$ kpc, however, as we see a flattening of the entropy profile for both SPG-hydro and MPG-hydro simulations. We speculate that this is due to the strong thermal feedback in our simulations within $r\simeq 1$ kpc. Our MHD runs also show similar behaviour to the hydro runs with partial thermal AGN feedback in the radial profile. The spherical heat injection by AGN feedback in addition to the thermal feedback due to Type Ia supernova feedback leads to overheating within the central $r\simeq 3$ kpc. Recent work by Mohapatra et al. (2025) suggests that modeling the Type Ia supernova feedback via discrete explosions rather than uniformly distributed heating helps to maintain the low central entropy within $r\simeq 5$ kpc although these simulations were run for shorter time ($\sim$ few 100 Myr) and lacked AGN feedback.
For our simulations, we have kept the thermal energy fraction at $25\%$ of the total injected energy for the AGN for all the runs except for MPG-hydro-kinetic run. This is because even tough AGN jets thermalise as they travel through the dense CGM, they still remain kinetically dominated at $r\sim 1$ kpc. Similar studies by Qiu et al. (2021) have equipartitioned the AGN energy in kinetic and thermal part, arguing the AGN jets thermalise close to SMBH through shock heating. Simulations of galaxy cluster halos by Li & Bryan (2014b, a) find that choice of AGN energy partition between kinetic and thermal fraction does not significantly affect the AGN-ICM interaction. However, we find that the partitioning of AGN energy in kinetic and thermal fraction has significant impact of AGN-CGM interaction. One of the possible reason might be that the galaxy clusters being larger halos have a much deeper potential well and higher ICM density and pressure compared to giant elliptical galaxies. Other possibility might be the way thermal energy feedback is implemented in our simulations. Li & Bryan (2014b) injects the thermal energy within the bipolar jet source region while in our simulations, AGN thermal energy is injected within a spherical volume within $r\sim 1$ kpc around the SMBH.
@@ Acknowledgements
The data underlying this article will be shared on reasonable request to the corresponding author.